What are the x intercept and the y intercept of the graph of y=1/3x -6
Answer by Alan3354(69443)

y=1/3x -6 is ambiguous.
==========
Is it y = (1/3)x -6 ?
Or y = 1/(3x) - 6 ?
Or y = 1/(3x -6) ?